# Knots and variance ordering of sequential Monte Carlo algorithms

### v1

@@ 5 Optimality of adapted knots
Adapted knots and knotsets, introduced in Examples 3.6 and 3.11 respectively, possess optimality properties that distinguish them from other knots. Adapted knots have the greatest variance reduction of any single knot, and $t$ applications of adapted knots will have the greatest variance reduction of any $t$ knots. Moreover, repeated application of adapted knots can reduce the variance to a fundamental value associated with importance sampling. This property indicates that the partial ordering by knots includes a model with optimal variance — indicating further suitability of knots as a tool for algorithm design. The first optimality result considers the application of a single knot or knotset.
Applying an adapted $t$-knot to a Feynman–Kac model results in the largest possible variance reduction of any single $t$-knot. Similarly, an adapted knotset will dominate any other knotset in terms of asymptotic variance reduction. This indicates that adapted knots should be appraised first before considering other types of knots that are compatible with the Feynman–Kac model at hand. The optimality of adapted knots is expressed formally in Theorem 5.1.
@@
Proof.
First consider the case of knots. For $t>0$ and $\mathcal{K}=(t,R,K)$, let $\mathcal{R}=(t,\mathrm{Id},R)$ then $\mathcal{R}\ast\mathcal{K}\ast\mathcal{M}=\mathcal{K}^{\diamond}\ast\mathcal{M}$ by Proposition A.4 and hence $\mathcal{K}^{\diamond}\ast\mathcal{M}\preccurlyeq\mathcal{K}\ast\mathcal{M}$. For a knot at time $t=0$, $\mathcal{K}=(0,R,K)$ where $R$ is a probability measure. As such, let $\mathcal{R}=(0,\delta_{0},K_{0})$ where $K_{0}(0,\cdot)=R$ to reach the same conclusion.
For knotsets, Proposition A.5 ensures the existence of a knot $\mathcal{R}$ such that $\mathcal{R}\ast\mathcal{K}\ast\mathcal{M}=\mathcal{K}^{\diamond}\ast\mathcal{M}$ for any model $\mathcal{M}$ and knotset $\mathcal{K}$. Therefore, $\mathcal{K}^{\diamond}\ast\mathcal{M}\preccurlyeq\mathcal{K}\ast\mathcal{M}$. ∎
Hence, in conjunction with Theorem 4.4, the asymptotic variance is lowest with adapted knots and knotsets. Further, from Proposition A.4 we can state that any sequence of $t$-knots applied to $\mathcal{M}$ is also dominated by the adapted $t$-knot applied to $\mathcal{M}$.
We can also deduce from Theorem 5.1 that the asymptotic variance can only be reduced beyond that of an adapted $t$-knot by using at least two non-trivial knots. Using the adapted $t$-knot followed by some other non-trivial knot at time $s\neq t$ would guarantee a further reduction in the variance for example.
Next we consider the case of multiple applications of knotsets by comparing models of the form $\mathcal{K}_{t}\ast\cdots\ast\mathcal{K}_{1}\ast\mathcal{M}$ to the sequence of adapted knotsets applied to $\mathcal{M}$.
@@
Let $t\in\mathbb{N}_{1}$ and consider two sequences of knotsets, $\mathcal{K}_{s}$ and $\mathcal{K}_{s}^{\star}$, over $s\in[0\,{:}\,t-1]$. Let $\mathcal{M}_{s}=\mathcal{K}_{s-1}\ast\mathcal{M}_{s-1}$ and $\mathcal{M}_{s}^{\star}=\mathcal{K}_{s-1}^{\star}\ast\mathcal{M}_{s-1}^{\star}$ for $s\in[1\,{:}\,t]$ and initial model $\mathcal{M}_{0}=\mathcal{M}_{0}^{\star}\in\mathscr{M}_{n}$. If $\mathcal{K}_{s}^{\star}$ is the adapted knotset for $\mathcal{M}_{s}^{\star}$ for all $s\in[0\,{:}\,t-1]$ then $\mathcal{M}_{t}^{\star}\preccurlyeq\mathcal{M}_{t}$.
Proof.
We have that $\mathcal{M}_{t}=\mathcal{K}_{t-1}\ast\mathcal{M}_{t-1}$ and hence, by Proposition A.5, there exists a knotset $\mathcal{R}_{t}$ that completes $\mathcal{K}_{t-1}$, that is $\mathcal{R}_{t}\ast\mathcal{M}_{t}=\mathcal{K}_{t-1}^{\diamond}\ast\mathcal{M}_{t-1}$ where $\mathcal{K}_{t-1}^{\diamond}$ is the adapted knotset for $\mathcal{M}_{t-1}$. We use Proposition A.7 to find
$\mathcal{R}_{t}\ast\mathcal{M}_{t}$ $=\mathcal{K}_{t-1}^{\diamond}\ast\mathcal{K}_{t-2}\ast\cdots\ast\mathcal{K}_{0}\ast\mathcal{M}_{0}$
$=\mathcal{J}_{t-1}\ast\cdots\ast\mathcal{J}_{1}\ast\mathcal{K}_{0}^{\star}\ast\mathcal{M}_{0}$
$=\mathcal{J}_{t-1}\ast\cdots\ast\mathcal{J}_{1}\ast\mathcal{M}_{1}^{\star},$
for some knotsets $\mathcal{J}_{s}$ for $s\in[1\,{:}\,t-1]$, where the first equality follows by definition of $\mathcal{M}_{t-1}$. The process of completing the first knotset (now $\mathcal{J}_{t-1}$) with a $\mathcal{R}_{t-1}$ and moving the adapted knot to the last position can be repeated until we have
$\mathcal{R}_{1}\ast\cdots\ast\mathcal{R}_{t-1}\ast\mathcal{R}_{t}\ast\mathcal{M}_{t}=\mathcal{M}^{\ast}_{t},$
and hence $\mathcal{M}_{t}^{\star}\preccurlyeq\mathcal{M}_{t}$. ∎
Theorem 5.2 states that $\mathcal{K}_{t-1}^{\star}\ast\cdots\mathcal{K}_{0}^{\star}\ast\mathcal{M}_{0}\preccurlyeq\mathcal{K}_{t-1}\ast\cdots\mathcal{K}_{0}\ast\mathcal{M}_{0}$ and hence a sequence of $t$ adapted knotsets have a greater variance reduction that any other sequence of $t$ knotsets.
As adapted knotsets reduce the variance in each application by the maximum amount of any knotset in each application, a natural question to ask is; to what extent can the asymptotic variance be reduced by repeated applications of adapted knotsets? Theorem 5.3 describes the minimal variance achievable by knotsets.
Theorem 5.3 (Minimal variance from knotsets).
@@
Proof.
$=\prod_{p=0}^{n-1}\eta_{p}(G_{p})\eta_{n}(G_{n}^{2})-\eta_{n}(\varphi)^{2}$
$=\gamma_{n}(1)\eta_{n}(G_{n}^{2})-\eta_{n}(\varphi)^{2}$
$=\gamma_{n}(G_{n}^{2})-\eta_{n}(\varphi)^{2}$
$=v_{n,n}(\varphi).$
@@
For the final model, $\mathcal{M}_{n}$, the only variance remaining in the model is at the terminal time. We can view the SMC algorithm on $\mathcal{M}_{n}$ with adaptive resampling as equivalent to an importance sampler using $\eta_{n}$ as the importance distribution and weight function $G_{n}(x_{n})\prod_{p=0}^{n-1}\eta_{p}(G_{p})=G_{n}(x_{n})\gamma_{n}(1)$. From the importance sampling view, we know that to reduce the remaining variance its minimal value, we will need to consider both the terminal potential and the test function of interest. Hence we note that terminal knots, introduced next, will need a treatment that reflect this, and is necessarily different from standard knots.

### v2

@@ 5 Optimality of adapted knots
Adapted knots and knotsets, introduced in Examples 3.6 and 3.11 respectively, possess optimality properties that distinguish them from other knots. Applying an adapted $t$-knot to a Feynman–Kac model results in the largest possible variance reduction of any single $t$-knot. Similarly, an adapted knotset will dominate any other knotset in terms of asymptotic variance reduction. This indicates that adapted knots should be appraised first before considering other types of knots that are compatible with the Feynman–Kac model at hand. The optimality of adapted knots and knotsets is expressed formally in Theorem 5.1.
@@
Hence, in conjunction with Theorem 4.4, the asymptotic variance is lowest with adapted knots and knotsets. Further, from Proposition B.4 we can state that any sequence of $t$-knots applied to $\mathcal{M}$ is also dominated by the adapted $t$-knot applied to $\mathcal{M}$.
We can also deduce from Theorem 5.1 that the asymptotic variance can only be reduced beyond that of an adapted $t$-knot by using at least two non-trivial knots. For example, using the adapted $t$-knot followed by some other non-trivial knot at time $s\neq t$ would guarantee a further reduction in the variance.
Next we consider the case of multiple applications of knotsets by comparing models of the form $\mathcal{K}_{m}\ast\cdots\ast\mathcal{K}_{1}\ast\mathcal{M}$ to the sequence of adapted knotsets applied to $\mathcal{M}$.
@@
Let $m\in\mathbb{N}_{1}$ and consider two sequences of knotsets, $\mathcal{K}_{t}$ and $\mathcal{K}_{t}^{\star}$, over $t\in[0\,{:}\,m-1]$. Let $\mathcal{M}_{t}=\mathcal{K}_{t-1}\ast\mathcal{M}_{t-1}$ and $\mathcal{M}_{t}^{\star}=\mathcal{K}_{t-1}^{\star}\ast\mathcal{M}_{t-1}^{\star}$ for $t\in[1\,{:}\,m]$ and initial model $\mathcal{M}_{0}=\mathcal{M}_{0}^{\star}\in\mathscr{M}_{n}$. If $\mathcal{K}_{t}^{\star}$ is the adapted knotset for $\mathcal{M}_{t}^{\star}$ for all $t\in[0\,{:}\,m-1]$ then $\mathcal{M}_{m}^{\star}\preccurlyeq\mathcal{M}_{m}$.
Theorem 5.2 states that if $\mathcal{K}_{t}^{\star}$ are adapted knotsets for $t\in[1\,{:}\,m]$, then $\mathcal{K}_{m}^{\star}\ast\cdots\ast\mathcal{K}_{1}^{\star}\ast\mathcal{M}\preccurlyeq\mathcal{K}_{m}\ast\cdots\ast\mathcal{K}_{1}\ast\mathcal{M}$, and hence a sequence of $m$ adapted knotsets have a greater variance reduction than any other sequence of $m$ knotsets, $\mathcal{K}_{m}$ for $t\in[1\,{:}\,m]$.
As adapted knotsets reduce the variance by the maximum amount of any knotset in each application, a natural question to ask is; to what extent can the asymptotic variance be reduced by repeated applications of adapted knotsets? Theorem 5.3 describes the minimal variance achievable by knots and/or knotsets.
Theorem 5.3 (Minimal variance from knots).
@@
For the final model, $\mathcal{M}_{n}$, the remaining variance is only incurred at the terminal time. We can view the SMC algorithm on $\mathcal{M}_{n}$ with adaptive resampling as equivalent to an importance sampler using $\eta_{n}$ as the importance distribution and weight function $G_{n}(x_{n})\prod_{p=0}^{n-1}\eta_{p}(G_{p})=G_{n}(x_{n})\gamma_{n}(1)$. From the importance sampling view, we know that to reduce the remaining variance its minimal value, we will need to consider both the terminal potential and the test function of interest. Hence we note that terminal knots, introduced next, will need a treatment that reflect this, and is necessarily different from standard knots.